# Reasoning Language Models: A Blueprint

### v1

@@ Abstract
Reasoning language models (RLMs), also known as Large Reasoning Models (LRMs), such as OpenAI’s o1 and o3, DeepSeek-V3, and Alibaba’s QwQ, have redefined AI’s problem-solving capabilities by extending large language models (LLMs) with advanced reasoning mechanisms. Yet, their high costs, proprietary nature, and complex architectures—uniquely combining Reinforcement Learning (RL), search heuristics, and LLMs—present accessibility and scalability challenges. To address these, we propose a comprehensive blueprint that organizes RLM components into a modular framework, based on a survey and analysis of all RLM works. This blueprint incorporates diverse reasoning structures (chains, trees, graphs, and nested forms), reasoning strategies (e.g., Monte Carlo Tree Search, Beam Search), RL concepts (policy, value models and others), and supervision schemes (Output-Based and Process-Based Supervision). We also provide detailed mathematical formulations and algorithmic specifications to simplify RLM implementation. By showing how schemes like LLaMA-Berry, QwQ, Journey Learning, and Graph of Thoughts fit as special cases, we demonstrate the blueprint’s versatility and unifying potential. To illustrate its utility, we introduce x1, a modular implementation for rapid RLM prototyping and experimentation. Using x1 and a literature review, we provide key insights, such as multi-phase training for policy and value models, and the importance of familiar training distributions. Finally, we outline how RLMs can integrate with a broader LLM ecosystem, including tools and databases. Our work demystifies RLM construction, democratizes advanced reasoning capabilities, and fosters innovation, aiming to mitigate the gap between “rich AI” and “poor AI” by lowering barriers to RLM development and experimentation.
@@
Survey of Reasoning Language Models, Survey of RLMs, Reasoning Language Model, Large Reasoning Model, RLM, LRM, Reasoning LLMs, Reinforcement Learning for LLMs, MCTS for LLMs, Large Language Model, LLM, Generative AI.
@@ 1 Introduction
The technical foundations of RLMs remain opaque and complex, compounding the accessibility challenge. Emerging analyses suggest that their design likely integrates elements such as Monte Carlo Tree Search (MCTS), reinforcement learning (RL), process-based supervision (PBS) [91, 55, 91, 55], and advanced in-context learning (ICL) techniques like Chain-of-Thought (CoT) [98] or Tree of Thoughts (ToT) [101], and possibly even retrieval-augmented generation (RAG) [52, 51, 37, 10].
@@
Our blueprint comprehensively encompasses the potential building blocks of RLMs, offering a flexible and modular framework. It incorporates a variety of reasoning structures, such as chains, trees, graphs, and even higher-order structures such as hierarchical (or nested) trees, along with numerous operations that transform and advance the reasoning process. The blueprint supports different granularities of reasoning steps, ranging from individual tokens to full sentences or structured segments. Additionally, it enables diverse training schemes, including Output-Based Supervision (OBS) and PBS. Furthermore, in order to illustrate the capability of the blueprint to accommodate novel design ideas, we describe several novel schemes and how they fit within the blueprint. One such example is Trace-Based Supervision (TBS), which extends PBS by incorporating labeled traces of traversal paths through entire reasoning structures, rather than just linear chains of reasoning steps. By unifying all these components, our blueprint serves as a versatile toolbox for constructing RLMs—ranging from simple models to sophisticated designs—tailored to specific reasoning tasks and performance objectives.
@@
We conduct a broad analysis of existing reasoning schemes (contribution #2), demonstrating how they fit into our blueprint as special cases. This analysis encompasses not only standard MCTS and reinforcement learning-based designs, such as LLaMA-Berry [105], but also models like QwQ [88]. Additionally, we include paradigms diverging from standard MCTS, such as Journey Learning [69], which redefines reasoning through implicit long-chain structures, and advanced structured prompting techniques like CoT [98], ToT [101], and Graph of Thoughts [6]. By mapping these diverse approaches to one blueprint, we showcase its versatility and expressive power, highlighting its ability to unify a wide range of reasoning methodologies within a coherent framework.
To demonstrate the utility of our framework, we introduce x1, a modular and user-friendly implementation [^2] https://github.com/spcl/x1 designed to simplify the process of developing and experimenting with new RLM architectures, covering not only training and inference, but also synthetic data generation (contribution #3). We design x1 to facilitate supporting various optimizations, design decisions, and overall scalability, such as batch processing, making it a well-suited foundation of experimentation infrastructure.
@@ 2.4.2 Explicit Reasoning Models
Figure 4: Overview of a general RLM design and core concepts. We provide a high-level overview (the top-left part), a more detailed medium-level overview (the top-right part), and a very detailed diagram showing the inference and training pipelines (the bottom part). A detailed specification of the inference pipeline can be found in Appendix C.1 and in Algorithm 1. Pipelines for different training phases and paradigms can be found in Appendix C.2–Appendix C.4, and in Algorithms 27. The data generation pipeline is detailed in Appendix D.
@@ 3.1.1 Inference
The details of how the structure evolves are usually governed by the MCTS scheme, enhanced with policy and value models (we also distinguish other reasoning strategies, described below). This approach, inspired by methods used in AlphaZero, ensures that the search process is both efficient and directed toward promising solutions. The policy model [Uncaptioned image] is responsible for generating new reasoning steps at each node, predicting the next most likely and logical steps to expand the reasoning process. Meanwhile, the value model [Uncaptioned image] evaluates the quality of a reasoning path starting at a given node, helping the system prioritize the most promising steps to follow. Sometimes, the reward model [^3] We use a naming scheme in which a model used to estimate the quality of a whole reasoning path starting at a given node, is called the value model, while a model used to estimate the quality of a given reasoning step, is called the reward model. could also be used, to assess the quality of an individual specific node and its corresponding reasoning step [Uncaptioned image]. In our blueprint, as detailed in the next section, we abstract the models into a more general notion of operators [Uncaptioned image] to enable more flexibility in how they are implemented.
@@ 3.1.2 Training
Training details depend on what model is trained (value, policy, reward, …). In general, we assume fine-tuning a model such as Llama. Here, we follow an approach where one first harnesses supervised data, usually coming from existing datasets such as PRM800K [55] [Uncaptioned image], which becomes a part of the framework supervised training data [Uncaptioned image] used in the supervised training pipeline [Uncaptioned image] to train some, or all, of the models considered in the blueprint [Uncaptioned image]. The second part of the overall training framework in RLMs is the unsupervised (self-learning) training pipeline, in which data is being continually generated and used to improve the models. The data can be obtained from inference [Uncaptioned image], assuming quality control [36], but also from a dedicated synthetic data generation pipeline that mirrors that of the inference [Uncaptioned image]. To collect the data, one executes the respective RLM pipeline for a given input task and gathers the results [Uncaptioned image]; depending on how detailed is the gathering process, the data collected can contain only output-based labels [Uncaptioned image], process-based labels [Uncaptioned image], or some other variant such as trace-based labels suggested in our blueprint, that generalize process-based samples to samples that contain also information about operators applied during the task solution process [Uncaptioned image]. All this data becomes a part of the replay buffer [Uncaptioned image] and is used in the unsupervised training scheme [Uncaptioned image] or it can also be used to train a model that would become an Implicit RLM [Uncaptioned image].
A detailed specification of the pipelines for different training phases and paradigms can be found in Appendix C.2, Appendix C.3, Appendix C.4, and in Algorithms 27. The data generation pipeline is detailed in Appendix D.
@@ 3.2 Encompassing Diverse RLM Architectures
In some RLM designs [105], a single node in the MCTS tree could represent an entire reasoning structure, such as a complete chain of reasoning steps. In this case, the action space involves transitioning between different reasoning chains rather than individual steps. This approach changes the nature of the search, as the focus shifts from iteratively constructing a single reasoning path to evaluating and refining entire structures within the search space. Our blueprint accommodates this with the concept of nesting, where a node in the reasoning structure can contain another reasoning structure.
@@ 4 Blueprint for Reasoning LMs
Figure 5: A blueprint for reasoning LMs. It consists of four main toolboxes: the reasoning scheme (the top part), operators (the bottom-left part), and models (the bottom-right part); pipelines are mentioned in the center and detailed in Appendix C.1 and in Algorithm 1 (the inference pipeline), Appendix C.2–Appendix C.4, and in Algorithms 27 (the training pipelines), and in Appendix D (the data generation pipeline).
@@ 4.2.3 Reasoning Strategy
Reasoning Strategy vs. Decoding Strategy It is crucial to distinguish reasoning strategies from token-level decoding strategies. While decoding strategies, such as greedy search and nucleus sampling [42], generate the internal token sequences within a reasoning step, reasoning strategies focus on the higher-level process of integrating and expanding reasoning steps within the reasoning structure.
@@ 4.4.2 Training Data Scope
The training data for RLMs can vary significantly in terms of how much of the reasoning structure it captures. We now outline two established approaches, output-based supervision (OBS) and process-based supervision (PBS). More details regarding both OBS and PBS can be found in Appendix B.1.
In output-based supervision (also known as a sparse training signal[25, 91] each training sample consists solely of the input and the corresponding output. For example, in mathematical problem-solving, a sample may include the task statement and the final solution, labeled as correct or incorrect. This approach is straightforward to implement, and the required data is relatively easy to collect. However, it can limit the model’s reasoning accuracy, as it provides minimal insight into the intermediate steps that led to the solution [55].
@@
OBS vs. PBS By varying the training data scope, developers can strike a balance between ease of data collection and the depth of reasoning insights provided to the model, with dense supervision generally offering improved performance at the cost of increased data complexity.
@@ 7.3 Models & Training Paradigms
Both the value and the policy model in x1 are fine-tuned versions of an LLM [^6] We currently use Llama-3.1-8B-Instruct as base model., without reliance on prompting, which is used in several other RLM architectures [36, 106]. This design decision aims to maximize the quality of results. We now outline briefly selected key aspects of how we train these models, full details can be found in Appendix B and C.3.
@@ 7.4 Enabling Scalability and Efficiency
7.5 Token Analysis
Our framework enables detailed analyses of strategies for generating and selecting fine-grained token-based reasoning steps. We consider variance, entropy, VarEntropy and the Gini coefficient as metrics for the analysis of the token probability distribution.
The variance in the probability distribution over the vocabulary could be a measure of uncertainty. A low variance would mean a high amount of similar values, indicating a flat distribution. However the variance does not capture well the form of the distribution.
Traditionally, entropy measures the uncertainty in a probability distribution. The information theoretic definition also describes the information content of a given systems state. A high entropy means a lot of information is necessary to quantify a system. This is the case if all positions have roughly the same probability, indicating a flat distribution. A distribution with high peaks would be characterized by a low entropy. The entropy is defined as
$H=-\sum_{i}p_{i}\log_{2}(p_{i})$
where $p_{i}$ is the probability of the $i$-th token.
VarEntropy is usually only applied in information theory and is defined as the variance of entropy, measuring its change. A high VarEntropy and a lower Entropy mean that a system has few very probable outcomes, while a low VarEntropy and high Entropy indicate a rather flat distribution. The VarEntropy is defined as
$\sum_{i}p_{i}(|\log(p_{i})|-|H|)^{2}$
The Gini coefficient is a traditional measure to gauge the inequality in a system, which plots the area covered by the sorted percentages of a group. A perfectly equal distribution, i.e. a flat distribution, would have a coefficient of 0. High inequality is characterized in turn by a value close to 1.

### v2

@@ Abstract
Reasoning language models (RLMs), also known as Large Reasoning Models (LRMs), such as OpenAI’s o1 and o3, DeepSeek-V3, and Alibaba’s QwQ, have redefined AI’s problem-solving capabilities by extending large language models (LLMs) with advanced reasoning mechanisms. Yet, their high costs, proprietary nature, and complex architectures—uniquely combining Reinforcement Learning (RL), search heuristics, and LLMs—present accessibility and scalability challenges. To address these, we propose a comprehensive blueprint that organizes RLM components into a modular framework, based on a survey and analysis of all RLM works. This blueprint incorporates diverse reasoning structures (chains, trees, graphs, and nested forms), reasoning strategies (e.g., Monte Carlo Tree Search, Beam Search), RL concepts (policy, value models and others), supervision schemes (Outcome-Based and Process-Based Supervision), and other related concepts (e.g., Test-Time Compute, Retrieval-Augmented Generation, agent tools). We also provide detailed mathematical formulations and algorithmic specifications to simplify RLM implementation. By showing how schemes like LLaMA-Berry, QwQ, Journey Learning, and Graph of Thoughts fit as special cases, we demonstrate the blueprint’s versatility and unifying potential. To illustrate its utility, we introduce x1, a modular implementation for rapid RLM prototyping and experimentation. Using x1 and a literature review, we provide key insights, such as multi-phase training for policy and value models, and the importance of familiar training distributions. Finally, we discuss scalable RLM cloud deployments and we outline how RLMs can integrate with a broader LLM ecosystem. Our work demystifies RLM construction, democratizes advanced reasoning capabilities, and fosters innovation, aiming to mitigate the gap between “rich AI” and “poor AI” by lowering barriers to RLM development and experimentation.
@@
Reasoning Language Model, Large Reasoning Model, Survey of Reasoning Language Models, Survey of RLMs, RLM, LRM, Reasoning LLMs, Reinforcement Learning for LLMs, MCTS for LLMs, Large Language Model, LLM, Generative AI.
@@ 1 Introduction
The technical foundations of RLMs remain opaque and complex, compounding the accessibility challenge. Emerging analyses suggest that their design likely integrates elements such as Monte Carlo Tree Search (MCTS) or Beam Search, reinforcement learning (RL), process-based supervision (PBS) [143, 83, 143, 83], and advanced in-context learning (ICL) techniques like Chain-of-Thought (CoT) [152] or Tree of Thoughts (ToT) [161], and possibly even retrieval-augmented generation (RAG) [79, 78, 53, 13].
@@
Our blueprint comprehensively encompasses the potential building blocks of RLMs, offering a flexible and modular framework. It incorporates a variety of reasoning structures, such as chains, trees, graphs, and even higher-order structures such as hierarchical (or nested) trees, along with numerous operations that transform and advance the reasoning process. The blueprint supports different granularities of reasoning steps, ranging from individual tokens to full sentences or structured segments. Additionally, it enables diverse training schemes, including Outcome-Based Supervision (OBS) and PBS, and the related Outcome & Process Reward Models (ORMs & PRMs). Next, in order to illustrate the capability of the blueprint to accommodate novel design ideas, we describe several novel schemes and how they fit within the blueprint. One such example is Trace-Based Supervision (TBS), which extends PBS by incorporating labeled traces of traversal paths through entire reasoning structures, rather than just linear chains of reasoning steps. By unifying all these components, our blueprint serves as a versatile toolbox for constructing RLMs—ranging from simple models to sophisticated designs—tailored to specific reasoning tasks and performance objectives.
@@
We conduct a broad analysis of existing reasoning schemes (contribution #2), demonstrating how they fit into our blueprint as special cases. This analysis encompasses not only standard MCTS and reinforcement learning-based designs, such as LLaMA-Berry [169], but also models like QwQ [140]. Additionally, we include paradigms diverging from standard MCTS, such as Journey Learning [113] or Beam Search, which redefines reasoning through implicit long-chain structures, and advanced structured prompting techniques like CoT [152], ToT [161], and Graph of Thoughts [9]. We also consider reasoning utilities such as Retrieval-Augmented Generation (RAG) and data stores, tools, and others. By mapping these diverse approaches to one blueprint, we showcase its versatility and expressive power, highlighting its ability to unify a wide range of reasoning methodologies within a coherent framework.
To demonstrate the utility of our framework, we introduce x1, a modular and user-friendly implementation [^2] https://github.com/spcl/x1 designed to simplify the process of developing and experimenting with new RLM architectures, covering not only training and inference, but also synthetic data generation (contribution #3). We design x1 to facilitate supporting various optimizations, design decisions, and overall scalability, such as batch processing, making it a well-suited foundation of experimentation infrastructure. We also discuss key aspects of deployment in cloud environments, ensuring that x1 can be seamlessly integrated into modern infrastructure for both research and production use cases.
@@ 2.4.2 Explicit Reasoning Models
Figure 4: Overview of a general RLM design and core concepts. We provide a high-level overview (the top-left part), a more detailed medium-level overview (the top-right part), and a very detailed diagram showing the inference and training pipelines (the bottom part). A detailed specification of the inference pipeline can be found in Appendix C.1 and in Algorithm 1. Details on the pipelines for different training phases and paradigms can be found in Appendices C.2 and C.3 as well as in Algorithms 27. The data generation pipeline is detailed in Appendix D.
@@ 3.1.1 Inference
The details of how the structure evolves are usually governed by the MCTS scheme, enhanced with policy and value models (we also distinguish other reasoning strategies, described below). This approach, inspired by methods used in AlphaZero, ensures that the search process is both efficient and directed toward promising solutions. The policy model [Uncaptioned image] is responsible for generating new reasoning steps at each node, predicting the next most likely and logical steps to expand the reasoning process. Meanwhile, the value model [Uncaptioned image] evaluates the quality of a reasoning path starting at a given node, helping the system prioritize the most promising steps to follow. Sometimes, a reward model [^3] We use a naming scheme in which a model used to estimate the quality of a whole reasoning path starting at a given node, is called the value model, while a model used to estimate the quality of a given reasoning step, is called the reward model. [Uncaptioned image] is used instead, to assess the quality of an individual specific node and its corresponding reasoning step. In our blueprint, as detailed in the next section, we abstract the models into a more general notion of operators [Uncaptioned image] to enable more flexibility in how they are implemented.
@@ 3.1.2 Training
Training details depend on what model is trained (value, policy, reward, …). In general, we assume fine-tuning a model such as Llama. Here, we follow an approach where one first harnesses supervised data, usually coming from existing datasets such as PRM800K [83] [Uncaptioned image], which becomes a part of the supervised training data [Uncaptioned image] used in the supervised training pipeline [Uncaptioned image] of the framework to train some, or all, of the models [Uncaptioned image] considered in the blueprint. The second part of the overall training framework in RLMs is the unsupervised (self-learning) training pipeline, in which training data is being continually generated [Uncaptioned image] and used to improve the models. The data can be obtained from inference, assuming quality control [52], but also from a dedicated synthetic data generation pipeline that mirrors that of the inference. To collect the data, one executes the respective RLM pipeline for a given input task and gathers the results [Uncaptioned image]; depending on how detailed the gathering process is, the data collected can contain only outcome-based labels [Uncaptioned image], process-based labels [Uncaptioned image], or some other variant such as trace-based labels [Uncaptioned image] suggested in our blueprint, that generalize process-based samples to samples that contain also information about operators applied during the task solution process. All this data becomes a part of the replay buffer [Uncaptioned image] and is used in the unsupervised training scheme [Uncaptioned image] or it can also be used to train [Uncaptioned image] a model that would become an Implicit RLM [Uncaptioned image].
A detailed specification of the pipelines for different training phases and paradigms can be found in Appendices C.2 and C.3 as well as in Algorithms 27. The data generation pipeline is detailed in Appendix D.
@@ 3.2 Encompassing Diverse RLM Architectures
In some RLM designs [169], a single node in the MCTS tree could represent an entire reasoning structure, such as a complete chain of reasoning steps. In this case, the action space involves transitioning between different reasoning structures rather than individual steps. This approach changes the nature of the search, as the focus shifts from iteratively constructing a single reasoning path to evaluating and refining entire structures within the search space. Our blueprint accommodates this with the concept of nesting, where a node in the reasoning structure can contain another reasoning structure.
@@ 4 Blueprint for Reasoning LMs
Figure 5: A blueprint for reasoning LMs. It consists of four main toolboxes: the reasoning scheme (the top part), operators (the bottom-left part), and models (the bottom-right part); pipelines are mentioned in the center and detailed in Appendix C.1 and in Algorithm 1 (the inference pipeline), Appendix C.2, Appendix C.3, and in Algorithms 27 (the training pipelines), and in Appendix D (the data generation pipeline).
@@ 4.2.3 Reasoning Strategy
Reasoning Strategy vs. Decoding Strategy. It is crucial to distinguish reasoning strategies from token-level decoding strategies. While decoding strategies, such as greedy search and nucleus sampling [60], generate the internal token sequences within a reasoning step, reasoning strategies focus on the higher-level process of integrating and expanding reasoning steps within the reasoning structure.
@@ 4.3.4 Evaluate Operators
4.3.5 Discussion: Test-Time Compute
One of the recent trends in next-generation LLMs [145, 95] is to shift from merely increasing model sizes to enhancing computational strategies during inference, a concept known as the test-time compute (TTC). This approach allocates additional computational resources during a model’s execution to improve performance, particularly in complex reasoning tasks. This methodology mirrors human cognitive processes, where increased deliberation is applied to more challenging problems.
Recent studies [131] indicate that optimizing test-time compute can be more effective than merely increasing model size. For instance, employing a compute-optimal strategy—where computational resources are adaptively allocated based on the problem’s complexity—can enhance efficiency by over four times compared to traditional methods. Moreover, in scenarios where smaller base models achieve moderate success rates, augmenting test-time compute enables them to outperform models up to 14 times larger.
While test-time compute offers significant benefits, it also presents challenges, related to – among others – resource allocation (determining the optimal amount of computational resources for each inference task requires sophisticated strategies to balance performance gains against computational costs), dynamic scaling (implementing adaptive compute strategies necessitates models capable of assessing problem difficulty in real-time and adjusting their computational efforts accordingly) [97], and hardware implications (the shift towards increased test-time computation may influence hardware requirements, putting more pressure on delivering specialized inference-focused hardware solutions).
Test-Time Compute in the Context of the Blueprint. Our blueprint offers mechanisms to dynamically allocate computational resources during inference to improve performance, particularly for more complex problems. By leveraging the modular structure of the blueprint, TTC can be effectively implemented through specific operators designed for reasoning tasks. We now provide several examples.
  • The Generate operator can be used to implement TTC by dynamically increasing the number of next reasoning steps generated for harder problems. For simpler tasks, the operator may only generate a minimal set of continuations. However, for more complex problems, the operator can be used to create a larger set of potential reasoning steps, thereby expanding the search space.
  • The Refine operator provides another avenue for implementing TTC by enhancing a given reasoning step multiple times for harder problems. In this approach, the operator iteratively improves the quality of a reasoning step, addressing ambiguities, rectifying errors, or improving clarity. For simpler tasks, the operator might only refine a step once, while for more complex reasoning, it can perform multiple enhancement iterations to ensure the output meets a higher standard of precision and robustness.
  • The Traversal operators, such as Select, enable the exploration of multiple reasoning paths at test time, offering another key mechanism for implementing TTC [171]. By using Select on several next reasoning steps, the model can dynamically expand its search tree for more challenging problems, thereby increasing the diversity and depth of reasoning paths under consideration. For example, in a complex task, the model might select multiple high-probability steps and explore their corresponding continuations in parallel. This approach facilitates broader exploration of the reasoning space, ensuring that promising paths are not prematurely discarded.
  • To efficiently manage the expanded set of possibilities, the blueprint allows integration with the Aggregate operator. This operator evaluates the generated reasoning paths and selects the most promising ones based on predefined criteria, such as the likelihood of correctness or the quality of intermediate steps. This combination ensures that while more computational resources are allocated for challenging tasks, only the most relevant paths are explored further, optimizing both accuracy and efficiency.
@@ 4.4 Models
Models are further categorized and discussed in detail in Appendix B; we discuss the variants of the value model (Q Value model, V Value model), we compare Process Reward and Outcome Reward models, and we formally identify a new variant of models, the Outcome-Driven Process Reward Model.
@@ 4.4.2 Training Data Scope
The training data for RLMs can vary significantly in terms of how much of the reasoning structure it captures. We now outline two established approaches, outcome-based supervision (OBS) and process-based supervision (PBS). More details regarding both OBS and PBS can be found in Appendix B.1.
In outcome-based supervision (also known as a sparse training signal[35, 143] each training sample consists solely of the input and the corresponding output. For example, in mathematical problem-solving, a sample may include the task statement and the final solution, labeled as correct or incorrect. This approach is straightforward to implement, and the required data is relatively easy to collect. However, it can limit the model’s reasoning accuracy, as it provides minimal insight into the intermediate steps that led to the solution [83].
@@
OBS vs. PBS By varying the training data scope, developers can strike a balance between ease of data collection and the depth of reasoning insights provided to the model, with dense supervision generally offering improved performance at the cost of increased data complexity. We detail these, and additional aspects of ORMs and PRMs in Pipelines for different training phases and paradigms can be found in Appendix B, Appendix C.2, Appendix C.3, and in Algorithms 27.
@@ 7.3 Models & Training Paradigms
Both the value and the policy model in x1 are fine-tuned versions of an LLM [^6] We currently use Llama-3.1-8B-Instruct as base model., without reliance on prompting, which is used in several other RLM architectures [52, 170]. This design decision aims to maximize the quality of results. We now outline briefly selected key aspects of how we train these models, full details can be found in Appendix B, C, and D.
@@ 7.4 Enabling Scalability and Efficiency
7.5 Blueprint for Efficient Scaling
Our blueprint can be deployed to AI HPC systems and clouds, as both systems provide the performance and resources necessary to scale RLMs. Deployment on HPC systems is straightforward: compute tasks are distributed across statically allocated nodes, connected with a low-latency and high-bandwidth interconnect, and with training data being available on a high-performance parallel filesystem. On the other hand, the cloud provides many configurable services that offer different trade-offs between performance, cost, and reliability. There, it becomes the user’s responsibility to choose the storage options and compute granularity that provides the best match for expected performance and cost. The architecture of our blueprint fits into the microservice architecture, with a clear separation of compute tasks, data storage, and coordination. This architecture helps to ease the configuration process, as different components of the system can be deployed, scaled, and optimized independently. In particular, the separation of value and policy servers allows them to be scaled separately according to the complexity of reasoning steps that might require different resource allocations to handle task generation and evaluation.
First, we outline the major decisions users must make before deploying the x1 scaling blueprint:
  • Deployment Training and inference tasks are typically allocated to virtual machines and containers, with the latter typically deployed as managed services with an orchestrator such as Kubernetes. There, x1 can benefit from modern frameworks like Ray [105] that hide the complexity of managing a service in a Kubernetes cluster.
  • Data Storage In the cloud, object storage provides automatic bandwidth scalability that allows scale computations operating on the same data. To overcome latency and power constraints, data can also be placed in in-memory caches like Redis and hybrid solutions that combine disks with flash memory [172].
  • Communication Requirements of the x1 blueprint differ from classical microservices, that rely on high-level abstractions like RPC and REST interfaces. RLM must utilize high-performance network fabrics offered by modern clouds, such as InfiniBand on Azure and Elastic Fabric Adapter (FBA) on AWS, both capable of achieving throughput of 400 Gb/s [39]. These are also available to training processes distributed across many GPUs, e.g., through specializations of the NVIDIA collectives library NCCL.
  • Parallelism We apply parallelism at multiple blueprint levels, including the classic data, model, and pipeline parallelism. These can scaled horizontally across a larger number of virtual machines and containers. On the other hand, reasoning steps can benefit from elastic scaling, like in distributed MCTS and Beam Search, where each path can be explored in parallel. There, containers can be allocated on the fly to support new paths and deallocated as soon as the parallelism scale of the computation decreases.
New developments in the machine learning infrastructure can significantly impact RLM deployment strategies:
  • Elastic Compute Computing tasks can be executed on ephemeral resources that trade the guaranteed lifetime and reliability for lower costs, such as spot virtual machines [101]. Serverless functions provide elasticity scalability with fine-grained pricing models [37], which can be a good fit for dynamically generated reasoning steps. However, serverless functions are stateless and suffer from cold starts, which requires optimization techniques dedicated to LLMs [47]. Furthermore, restricted network communication in functions forces the adoption of new communication protocols [70, 36].
  • GPU Management Cloud rental of GPU devices is particularly expensive, and procuring a sufficient number of devices can be challenging, specifically when constrained to a single cloud region. Given the large compute and memory requirements of base models, space-sharing might not be feasible. On the other hand, time-sharing of GPU devices between different x1 services could be a viable alternative, but it is currently constrained by large memory allocations and the cost of swapping model checkpoints between CPU and GPU memory. To increase resource utilization, new techniques for efficient GPU checkpoint and restore are needed [47].
  • Parameter-Efficient Resource Sharing Resource-sharing can be further enhanced by utilizing a shared base model architecture for the policy and value models, while dynamically swapping task-specific parameter layers - such as Low-Rank Adaptation [62], prefix tuning [81], or other adapter layers - on the GPU during inference. These modular strategies keep the base model loaded in device memory and replace only the lightweight task-specific layers, eliminating redundant loading and reducing both latency and memory usage. An example of an RLM, which uses a shared base model with separate additional linear layers for policy and value model, is AlphaMath [23].
  • Cross-Region Deployment Cloud applications are often deployed in a single region to avoid the performance and cost of cross-region data access. However, workloads can be scheduled globally, suspended, and migrated across regions to avoid hardware resource exhaustion and achieve lower carbon emissions [33, 153].
7.6 Example Analysis: Token Probability Distributions
(a) 1st example
(b) 2nd example
(c) 3rd example
(d) 4th example
Figure 7: Four examples of model output with highlighted tokens indicating uncertainty levels. The outputs have been color-coded to reflect the confidence levels of the model’s token predictions. Tokens are highlighted in purple when the highest probability is below 0.8 (indicating lower certainty without significant contention), in blue when the second-highest probability exceeds 0.1 (indicating contention, where another token is a close alternative), and in red when both conditions are met (indicating high uncertainty). These examples illustrate varying levels of prediction confidence and contention in reasoning steps, emphasizing regions of high ambiguity or competition between plausible continuations. This type of visual analysis is useful for identifying points in the reasoning process where the model lacks confidence or is torn between alternatives, guiding refinements in reasoning strategies and model design. It also helps pinpoint critical areas where additional supervision or context may improve model performance.
As an illustrative example, we use the framework to directly leverage the token probability distribution, thereby facilitating the use of associated properties—such as entropy and variance—for guiding subsequent reasoning decisions. By focusing on these probabilistic characteristics, the framework can help identify when to expand a given reasoning step. Using token probability distributions can be used for navigating the reasoning based on both coarse and fine steps. To support this analysis, the x1 implementation includes scripts that provide insights into token-level metrics, such as entropy fluctuations and distribution patterns, to inform reasoning strategies.
7.6.1 Relevance of Token Probability Distribution
The token probability distribution provides critical information about the likelihood of different next-step candidates in a reasoning process. By examining this distribution, we can gain insight into how certain tokens dominate or diversify the reasoning space, and in turn, guide more informed decisions about which step to take next.
We now list a few scenarios where different token distributions offer insights into which reasoning decision is best to take at a given step.
  • Flat Token Distribution. A flat probability distribution occurs when all tokens have roughly equal probabilities. In this scenario, there is significant uncertainty about which step is the best to choose because no single token stands out as a clear candidate. This can make the reasoning process more exploratory, as the model may need to consider multiple tokens equally and rely on additional strategies—such as external heuristics or learned policies—to identify the most promising step. While this can foster exploration, it may also lead to inefficiencies since the model might need to evaluate many equally plausible paths before finding an optimal solution. Another decision that could be taken in such a scenario, is to delay initiating a reasoning step till the token distribution changes to be more skewed.
  • Skewed Distribution with One Dominant Token. When one token has a much higher probability than others, the distribution is highly skewed. This often signals that the model is confident about the next step in the reasoning process. If the dominant token corresponds to a logical or well-supported continuation, this confidence can streamline decision-making and reduce computational overhead. However, if the model’s confidence is misplaced—perhaps due to biases in the training data or a lack of context—relying on a single dominant token may cause the reasoning process to follow a suboptimal path. In such cases, it’s crucial to assess whether the high-probability token genuinely represents the most logical next step or if additional validation is needed.
  • Skewed Distribution with Multiple High-Probability Tokens. In some cases, the distribution may be skewed with a small set of tokens receiving much higher probabilities than others. This indicates that the model sees several plausible continuations, each with a reasonable chance of being correct. While this is generally a positive sign—offering a diversity of credible options—it also complicates the decision-making process. The reasoning strategy must weigh the trade-offs between these top candidates, considering not only their individual probabilities but also how each choice impacts the subsequent reasoning trajectory. This scenario highlights the need for effective evaluation metrics (like entropy or Gini coefficient) to help select the step that contributes most to reaching the correct or desired outcome.
By analyzing token probability distribution and identifying the cases above and others, reasoning strategies can, for example, improve efficiency (identifying when a distribution is flat allows the reasoning algorithm to focus on diversification or introduce additional constraints to narrow down choices), enhance decision confidence (recognizing when one token is dominant can help expedite decisions, provided the model’s confidence is well-founded), or foster balanced exploration (detecting multiple high-probability tokens facilitates exploring various credible paths without being overly committed to a single option).
7.6.2 Analyzing Token Probability Distribution
To understand the form of a token probability distribution, we examine variance, entropy, VarEntropy, and the Gini coefficient as key metrics that offer distinct perspectives on the distribution’s shape and characteristics.
Variance provides a broad measure of uncertainty by reflecting how spread out the probabilities are across the vocabulary. When variance is low, the probabilities are nearly uniform, indicating a flat distribution. However, variance alone does not capture the specific structure or shape of the distribution. For example, two distributions can have the same variance but differ in their overall form, such as one having multiple minor peaks versus another being nearly uniform with a single dominant token. To address this, we consider further measures below.
Entropy has long been a standard measure of uncertainty and information content in a probability distribution. Higher entropy corresponds to greater unpredictability—requiring more information to describe the system’s state. For instance, if all tokens have nearly equal probabilities, the entropy is high, reflecting a flat distribution. In contrast, low entropy occurs when a small number of tokens dominate, resulting in a skewed distribution. The entropy of a distribution is given by $H=-\sum_{i}p_{i}\log_{2}(p_{i}),$ where $p_{i}$ is the probability of the $i$-th token. This metric provides valuable insight into whether the distribution is diffuse and exploratory or concentrated and decisive.
VarEntropy extends this analysis by measuring the variability of entropy itself, thus offering a dynamic view of how uncertainty changes. A high VarEntropy combined with low entropy often indicates a sharp, focused distribution with a few dominant outcomes. Conversely, low VarEntropy and high entropy typically reflect a flat, uniform distribution where no single token stands out. The VarEntropy is defined as $\sum_{i}p_{i}(|\log(p_{i})|-|H|)^{2}$. This metric captures the nuanced shifts in distribution shape, helping to pinpoint how tightly probabilities cluster around certain tokens versus how broadly they spread.
The Gini Coefficient, traditionally used to measure inequality, provides another lens on the form of the distribution. A perfectly equal distribution has a Gini coefficient of 0, signifying that all tokens have identical probabilities. A Gini coefficient closer to 1 indicates high inequality, where a few tokens hold most of the probability mass. By visualizing the cumulative distribution of sorted probabilities, the Gini coefficient highlights how the probability is concentrated or dispersed.
Together, these metrics—variance, entropy, VarEntropy, and Gini—enable a detailed examination of token probability distributions. By leveraging each metric’s unique strengths, we can effectively characterize whether a distribution is flat, skewed with a dominant token, or skewed across several highly probable tokens, ultimately guiding more informed decisions in reasoning and model development.